# End-to-End Crystal Structure Prediction from Powder X-Ray Diffraction

### v2

@@ 2.4 XtalNet Can Generate Crystal Structure Conditioned on PXRD
In order to generate superior corresponding crystal structure from PXRD, we can generate multiple candidate crystal structures and give these candidates reasonable ranks. As the diffusion generation process includes randomness, more candidates indicate more chance to include precise crystal structure. Here we use CPCP module mentioned above to calculate the similarity score between the target PXRD pattern and the generated crystal structures, thereby enabling ranking. As shown in Fig.3a, the match rates of hMOF-100 and hMOF-400 both increase as the number of candidates increases, which implies the multiple generations are useful for obtaining precise structure.
@@
Figure 3: Performance of XtalNet in Generating Crystal Structures. a, the match rates for hMOF-100 and hMOF-400 datasets with different number of top rank generated crystal structure candidates. b,c,the RMSE statistics for hMOF-100 and hMOF-400 dataset, indicating that XtalNet can generate highly accurate crystal structures from PXRD data for a significant proportion of cases. ]d, visual comparison of generated crystal structures and their simulated PXRD patterns against ground truth, highlighting the model’s performance in generating metal-connecting parts of MOFs and maintaining high similarity in PXRD patterns.
@@
The RMSE statistics of two dataset is shown in Fig. 3b and c. Here we use RMSE of best generation results of each test sample as the statics data. The cases of RMSE lower than 0.05 occupy quit a bit proportion comparing with other RMSE range cases, which means XtalNet can generate very precise crystal structure from PXRD for these cases. At the same time, the proportion of samples’ RMSE lower than 0.5 accounts for more than half of the total, while the RMSE of 0.5 represents the results achieves acceptable level.
@@
2.4.1 Effective Integration of PXRD Information into Crystal Generation Process by CCSG Module
Incorporating PXRD features into the crystal structure generation process is non-trivial, as there are multiple potential designs that can result in varying task performance levels. As illustrated in Fig. 3e, we compare the match rate of different designs in hMOF-100 dataset, as the match rate is a representative metric of the generation task. The Feat. node denotes that PXRD features are added as a new node in the crystal structure network, while Feat. cat indicates that PXRD features are concatenated with all crystal nodes. The P symbolizes that the PXRD feature extractor of the CCSG module is initialized from the CPCP module pre-training, and F signifies that the PXRD feature extractor of the CCSG module is frozen. Our strategy achieves the best results in comparison.
@@ 2.4.2 Hierarchical Optimization in the CCSG Module Generation Process
The generation process in our CCSG module is derived from the denoising process, which can be interpreted as a continuous update of atomic positions and lattice matrix. To demonstrate the interpretability of the denoising process, we visualize the crystal structure of the same sample at different diffusion time steps in Fig. 3f.
@@ 2.4.3 Evaluation of XtalNet in Diverse System Sizes and Elemental Compositions
XtalNet can be applied to systems with varying numbers of atoms in the unit cell, effectively accommodating diverse system sizes. The match rate and structure number corresponding to different system sizes in the training set are depicted in Fig. 4a, alongside visualizations of various system sizes. The RMSE of best generation result for different system sizes is illustrated in Fig. 4b. Here All results are derived in hMOF-400 dataset. As the system size increases, the match rate decreases, and the RMSE also escalates. This trend can be attributed to two factors: the increasing complexity of structure generation as the system size expands, and the reduced number of larger system structures, potentially leading to imbalanced and insufficient training.
XtalNet is also applicable to systems containing different metal elements. The hMOF-400 dataset primarily comprises four metal element types: V, Cu, Zn and Zr. The match rates and RMSE for structures containing distinct metal elements are displayed in Fig. 4c and d. It can be observed that the RMSE increases with the average number of atoms in the unit cell, indicating that more complex crystal structures yield poorer prediction results, which aligns with general expectations. Notably, the match rate for crystals containing Zn is significantly higher than those with Cu, which may be due to the larger number of samples containing Zn. An increased sample size can result in more comprehensive training and optimization. Consequently, the performance is influenced by both the number of samples and the complexity of the system.
@@ 3 Discussion
The development and implementation of XtalNet represent a significant leap forward in the field of both crystal structure prediction and automated powder X-ray diffraction (PXRD) data analysis. Our model, which employs an end-to-end deep learning framework, has demonstrated the capability to accurately predict crystal structures that match given PXRD patterns without reliance on external databases. This is a substantial departure from traditional methods that often require extensive manual intervention and database matching, which can be time-consuming and yield suboptimal results due to the incompleteness of database coverage.
The success of XtalNet in generating crystal structures, particularly those with high atom numbers in the unit cell, underscores the model’s ability to handle complex organic systems such as metal-organic frameworks (MOFs). The contrastive learning approach and diffusion-based conditional generation used in XtalNet have proven effective in establishing a one-to-one mapping between the PXRD space and crystal structure space, thereby reducing ambiguity in the prediction of multiple stable crystal structures for a given chemical composition.
The results obtained from our curated datasets, hMOF-100 and hMOF-400, are particularly promising. The high top-10 hit ratio and match rate on the database retrieval and structure generation tasks indicate that XtalNet can effectively retrieve and generate crystal structures that closely match PXRD patterns. This level of accuracy is crucial for applications in material science, where precise structural information is essential for understanding material properties and guiding the design of new materials with tailored characteristics.
However, despite these advancements, there are limitations to the current study that warrant further exploration. While XtalNet has demonstrated impressive results in generating crystal structures from simulated PXRD data, the application of this model to real experimental PXRD data presents a unique set of challenges. The complexity and variability inherent in experimental data, such as noise, peak broadening, and preferred orientation effects, can significantly impact the accuracy and reliability of the crystal structure predictions made by XtalNet.
In conclusion, XtalNet represents a significant advancement in the automation of PXRD analysis and crystal structure prediction. The model’s ability to integrate experimental characterization with crystal structure prediction offers a more efficient and accurate approach to material characterization. While there are challenges ahead, the potential of this technology to revolutionize the field of materials science is immense. Future research should focus on expanding the scope of XtalNet, improving its ability on real experimental PXRD data, and exploring its integration with other experimental techniques to further enhance the predictive modeling of materials.
@@ 4 Methods
5 Methods
The unit cell constitutes the fundamental repeating unit that serves as the foundational building block for characterizing 3D crystal structures. A unit cell can be defined as ${\mathcal{M}}=({\bm{A}},{\bm{F}},{\bm{L}})$, where ${\bm{A}}\in{\mathbb{R}}^{N}$ denotes atom types, ${\bm{F}}\in[0,1)^{3\times N}$ encompasses fractional coordinates of the atoms, ${\bm{L}}\in{\mathbb{R}}^{3\times 3}$ represents the lattice matrix containing three basic vectors to delineate the periodicity of the crystal, and $N$ signifies the number of atoms in each unit cell. Given a PXRD pattern ${\bm{C}}_{XRD}$, based on prior knowledge, we can presume that both ${\bm{A}}$ and $N$ are provided. Consequently, conditional CSP aims to learn the conditional distribution $p({\bm{L}},{\bm{F}}\mid{\bm{A}},{\bm{C}}_{XRD})$.
Our approach integrates two specialized neural networks to address dual tasks involving PXRD and crystal structure data. We employ a PXRD Feature Extractor for obtaining crucial PXRD features and a Crystal Structure Network for encoding and update crystal structures, encompassing aspects such as lattice parameters and fractional atomic positions. The first task is to align PXRD and crystal structure data utilizing a contrastive learning approach, analogous to CLIP [22]’s method for images and text. The second task concentrates on generating crystal structures based on features predicted by the PXRD Feature Extractor.
@@ 5.1 Model Achitecture
In our research, we adapt a Modified Equivariant Graph Neural Network (EGNN), called CSPNet [4], from the DiffCSP framework to serve as the Crystal Structure Network ($f_{CSP}$), leveraging its proven competence in processing crystallographic data. To align the model with our unique research objectives, we have instituted several alterations. Originally, the CSPNet constructed node representations by amalgamating node and temporal features, subsequently undergoing a linear transformation to generate the ultimate node features. Pertinent to the diffusion process, we have incorporated an additional supervisory signal absent in the original CSPNet. This is achieved by fusing the node features with both temporal and condition-specific PXRD features before subjecting them to a linear layer, thus deriving the final node representations.
@@
${\bm{I}}_{node}=\left\{\begin{aligned} \text{Linear}({\bm{P}},t,\text{Embeddin}({\bm{A}})),\text{in CCSG process}\\ \text{Embeddin}({\bm{A}}),\text{in CPCP process}\end{aligned}\right.$
${\bm{C}},\hat{\varepsilon}_{{\bm{F}}},\hat{\varepsilon}_{{\bm{L}}}=f_{CSP}({\bm{I}}_{node},{\bm{F}},{\bm{L}}),$
Where ${\bm{P}}$ denotes PXRD feature, $t$ denotes time embedding, ${\bm{C}}$ denotes crystal feature, $\varepsilon_{{\bm{F}}}$ denoising term on fractional coordinate and $\varepsilon_{{\bm{L}}}$ denotes denoising term on lattice, while $P$ and $t$ is the optional inputs. Within our contrastive learning paradigm, we employ the mean of final all node features to obtain crystal feature ${\bm{C}}$, which encapsulates the aggregate structural characteristics.
@@
The pretrained PXRD feature extractor imparts conditional PXRD information that steer the direction of the crystal structure denoising. The PXRD features serve as a guide throughout this procedure, ensuring the synthesized crystal structure is not only in agreement with the PXRD data but also conforms to the material system’s intrinsic physical and chemical properties. This strategic enhancement integrates critical supervisory data into the model’s architecture.
5.2 Loss Functions
Contrastive Learning Loss
We propose a contrastive learning framework designed to synergize two heterogeneous data types: PXRD patterns and crystal structures. This framework draws inspiration from the Contrastive Language-Image Pretraining (CLIP) approach, which effectively bridges the gap between visual and textual information. Our aim is to construct a joint embedding space in which the representations of PXRD and crystal data are cohesively aligned. This facilitates cross-modal correlations and bolsters the feature learning process. To achieve this, we employ the InfoNCE loss within each batch to attract paired modal data points while repulsing unpaired ones. The intended outcome is to obtain embeddings for both modalities that can accurately encapsulate the characteristics of the objects.
In the context of contrastive learning, the InfoNCE loss is defined as follows, $\tau$ is the temperature coefficient used to scale the similarity computation, ${\bm{P}}_{i}$ is the $i$th PXRD feature and ${\bm{C}}_{i}$ is the $i$th crystal structure feature.

### v3

@@ 2.4 XtalNet Can Predict Crystal Structure Conditioned on PXRD
In pursuit of generating superior corresponding crystal structures from PXRD, it is advantageous to produce a multitude of candidate structures, subsequently assigning them reasonable ranks. As the diffusion generation process includes randomness, more candidates indicate more chance to include precise crystal structure. In this study, we generate 20 crystal structure candidates for the subsequent ranking process. Utilizing the previously mentioned CPCP module, we calculate the similarity score between the target PXRD pattern and the generated crystal structures, thereby facilitating their ranking. The match rate is determined among the top-k ranking candidates, while the RMSE corresponds to the best candidate within the top-k ranking group.
@@
Figure 3: Performance of XtalNet in Crystal Structure Prediction. a, the match rates for hMOF-100 and hMOF-400 datasets with different number of top rank generated crystal structure candidates. b,c,the RMSE statistics for hMOF-100 and hMOF-400 dataset, indicating that XtalNet can generate highly accurate crystal structures from PXRD data for a significant proportion of cases. d, visual comparison of generated crystal structures and their simulated PXRD patterns against ground truth, highlighting the model’s performance in generating metal-connecting parts of MOFs and maintaining high similarity in PXRD patterns. e, performance of different architectures, demonstraing the reasonableness of XtalNet. Feat. node denotes PXRD feature is added as a new node, Feat. cat denotes PXRD feature is concated with original node features, P denotes PXRD feature extractor is pretrained by CPCP module and F denotes PXRD feature extractor is frozen during CCSG training. f, diffusion trajectory of generating a crystal structure, showing the interpretability of the denoising process.
The assessment of the hMOF-100 dataset is performed on the entire test set, while the evaluation of the hMOF-400 dataset is conducted on 3220 samples, randomly chosen from the test set, owing to the computational resources constraint. As shown in Figure 3a, the match rates of hMOF-100 and hMOF-400 both increase as the number of candidates increases, which implies the multiple generations are useful for obtaining precise structure. The top-10 match rate of hMOF-100 dataset and hMOF-400 dataset also achieves 90.2% and 79% respectively.
@@
The RMSE statistics of two dataset is shown in Figure 3b and c. Here we use RMSE of best generation results of each test sample as the statics data. The cases of RMSE lower than 0.05 occupy quit a bit proportion comparing with other RMSE range cases, which means XtalNet can generate very precise crystal structure from PXRD for these cases. At the same time, the proportion of samples’ RMSE lower than 0.5 accounts for more than half of the total, while the RMSE of 0.5 represents the results achieves acceptable level.
@@
2.4.1 Effective Integration of PXRD in Crystal Generation
Incorporating PXRD features into the crystal structure generation process is non-trivial, as there are multiple potential designs that can result in varying task performance levels. As illustrated in Figure 3e, we compare the match rate of different designs in hMOF-100 dataset, as the match rate is a representative metric of the generation task. Here we randomly select 1000 samples from test set to evaluate the performance. The Feat. node denotes that PXRD features are added as a new node in the crystal structure network, while Feat. cat indicates that PXRD features are concatenated with all crystal nodes. The P symbolizes that the PXRD feature extractor of the CCSG module is initialized from the CPCP module pre-training, and F signifies that the PXRD feature extractor of the CCSG module is frozen. Our strategy achieves the best results in comparison.
@@ 2.4.2 Hierarchical Optimization in the CCSG Module Generation Process
The generation process in our CCSG module is derived from the denoising process, which can be interpreted as a continuous update of atomic positions and lattice matrix. To demonstrate the interpretability of the denoising process, we visualize the crystal structure of the same sample at different diffusion time steps in Figure 3f.
@@ 2.4.3 Evaluation of XtalNet in Diverse System Sizes and Elemental Compositions
XtalNet can be applied to systems with varying numbers of atoms in the unit cell, effectively accommodating diverse system sizes. The match rate and structure number corresponding to different system sizes in the training set are depicted in Figure 4a, alongside visualizations of various system sizes. The RMSE of best generation result for different system sizes is illustrated in Figure 4b. Here All results are derived in hMOF-400 dataset. As the system size increases, the match rate decreases, and the RMSE also escalates. This trend can be attributed to two factors: the increasing complexity of structure generation as the system size expands, and the reduced number of larger system structures, potentially leading to imbalanced and insufficient training.
XtalNet is also applicable to systems containing different metal elements. The hMOF-400 dataset primarily comprises four metal element types: V, Cu, Zn and Zr. The match rates and RMSE for structures containing distinct metal elements are displayed in Figure 4c and d. It can be observed that the RMSE increases with the average number of atoms in the unit cell, indicating that more complex crystal structures yield poorer prediction results, which aligns with general expectations. Notably, the match rate for crystals containing Zn is significantly higher than those with Cu, which may be due to the larger number of samples containing Zn. An increased sample size can result in more comprehensive training and optimization. Consequently, the performance is influenced by both the number of samples and the complexity of the system.
@@ 3 Discussion
The development and implementation of XtalNet represent a significant leap forward in the field of both crystal structure prediction and automated PXRD data analysis. Our model, which employs an end-to-end deep learning framework, has demonstrated the capability to accurately predict crystal structures that match given PXRD patterns without reliance on external databases. This is a substantial departure from traditional methods that often require extensive manual intervention and database matching, which can be time-consuming and yield suboptimal results due to the incompleteness of database coverage.
XtalNet is a method suitable for predicting any crystal structure system, and it does not impose any requirements on the system itself. Therefore, as long as there is a corresponding reasonable dataset, it can be used for both organic and inorganic systems. Here, we chose the more challenging organic system, MOF, to validate our method. The success of XtalNet in generating crystal structures underscores the model’s ability to handle complex systems such as metal-organic frameworks (MOFs). The contrastive learning approach and diffusion-based conditional generation used in XtalNet have proven effective in establishing a one-to-one mapping between the PXRD space and crystal structure space, thereby reducing ambiguity in the prediction of multiple stable crystal structures for a given chemical composition.
Due to the challenges associated with collecting PXRD experimental data, there is indeed a scarcity of such data. To address this limitation, we utilized GSAS software to simulate the PXRD data, generating a relatively large dataset for training purposes. This approach effectively mitigated the impact of data scarcity and resulted in promising outcomes. The high top-10 hit ratio and match rate indicate that XtalNet can effectively retrieve and generate crystal structures that closely match PXRD patterns, which is crucial for applications in material science as precise structural information is essential for understanding material properties and guiding the design of new materials with tailored characteristics.
However, despite these advancements, there are some limitations to the current study that warrant further exploration. XtalNet has demonstrated impressive results in generating crystal structures from simulated PXRD data, but the application of this model to real experimental PXRD data presents a unique set of challenges. The complexity and variability inherent in experimental data, such as noise, peak broadening, effect of solvent and preferred orientation, can significantly impact the accuracy and reliability of the crystal structure predictions made by XtalNet. Consequently, the application of deep learning techniques under complex experimental conditions remains a challenge. In the future, more data augmentation and simulation data enhancement can be taken to improve the performance.
@@ 4 Methods
The unit cell constitutes the fundamental repeating unit that serves as the foundational building block for characterizing 3D crystal structures. A unit cell can be defined as ${\mathcal{M}}=({\bm{A}},{\bm{F}},{\bm{L}})$, where ${\bm{A}}\in{\mathbb{R}}^{N}$ represents the atom types, ${\bm{F}}\in[0,1)^{3\times N}$ denotes the fractional coordinates of the atoms within the unit cell, and ${\bm{L}}\in{\mathbb{R}}^{3\times 3}$ is the lattice matrix representing the periodicity of the crystal, defined by three basis vectors derived using the Niggli algorithm [46], and $N$ signifies the number of atoms in each unit cell.For practical crystal identification problems based on PXRD, we can assume that the PXRD pattern ${\bm{C}}_{XRD}$, the atom types ${\bm{A}}$, and the number of atoms $N$ are provided. In our setup, conditional crystal structure prediction (CSP) aims to model the conditional distribution $p({\bm{L}},{\bm{F}}\mid{\bm{A}},{\bm{C}}_{XRD})$, learning the relationship between the crystal lattice, fractional coordinates, atom types, and PXRD patterns to predict the most probable crystal structure.
In our approach, we utilize paired PXRD and crystal structure data during the training phase. This method incorporates two specialized neural networks: a PXRD Feature Extractor, which derives essential features from PXRD patterns, and a Crystal Structure Network, which extracts structural features from given crystal structures and predicts their subsequent states. To effectively train these networks, we adopt a dual-task framework. The first task involves aligning the PXRD features produced by the PXRD Feature Extractor with the crystal structure features obtained from the Crystal Structure Network. This task employs a contrastive learning strategy, akin to the approach utilized by CLIP [35], aimed at aligning data across different modalities. The second task focuses on generating crystal structures using the Crystal Structure Network, based on features predicted by the PXRD Feature Extractor.
@@ 4.1 Model Achitecture
To identify patterns in Powder X-ray Diffraction (PXRD) data, we developed a neural network called the PXRD Feature Extractor ($f_{PXRD}$), which is based on the BERT architecture [39]. PXRD data are represented as curves with diffraction angles on the x-axis and intensities on the y-axis. Before inputting the PXRD data into the model, we preprocess the data by normalizing the intensities and smoothing the curves to reduce noise. This step ensures that the input data is standardized, facilitating better learning and reducing the impact of experimental variability.
@@
To preprocess the PXRD data for model input, ${\bm{A}}_{xrd}$ is treated as discrete positions indicative of peak locations, which are converted into positional embeddings using an embedding layer, similar to the position encodings in traditional Transformer models. Since ${\bm{I}}_{xrd}$ represents continuous intensity values, we use a multilayer perceptron (MLP) network to transform these values into input embeddings. Additionally, to enhance the representation capability of the model, a unique trainable embedding is added at the beginning of the tokenized sequence, similar to the ”[CLS]” token in BERT models. This embedding serves as a summary representation, or ”header,” for the PXRD data. The PXRD header feature, ${\bm{P}}=f_{PXRD}({\bm{A}}_{xrd},{\bm{I}}_{xrd})$, is subsequently used as the comprehensive representation of the entire PXRD data for contrastive pretraining and as a conditioning input for the diffusion models. The PXRD feature extractor is pre-trained in the Contrastive PXRD-Crystal Pretraining (CPCP) module to align PXRD features ${\bm{P}}$ with crystal structure feature ${\bm{C}}$. During the training of the Conditional Crystal Structure Generation (CCSG) module, the PXRD feature extractor is initialized with pre-trained parameters and remains frozen to maintain the learned feature representations.
@@
In our study, we utilize a Modified Equivariant Graph Neural Network (EGNN), referred to as CSPNet [21], derived from the DiffCSP framework, to serve as the Crystal Structure Network ($f_{CSP}$). This network is specifically tailored for handling crystallographic data. To align with our specific research goals, several modifications were made to the original CSPNet. For the diffusion process, CSPNet constructs node features by initially combining atom attributes (such as atom types) with time embedding. These combined features are then transformed to produce final node representations through Message Passing Neural Networks (MPNN). Additionally, we introduce a novel conditioning signal, incorporating temporal embeddings and condition-specific PXRD features ${\bm{P}}$, to better guide the denoising process. This is achieved by integrating node features with both time embedding and condition-specific PXRD features ${\bm{P}}$ before passing them through a linear transformation layer:
@@
${\bm{N}}^{0}=\left\{\begin{aligned} \text{Linear}(\text{cat}({\bm{P}},\text{Embedding}({\bm{A}}),t)),\text{in CCSG module}\\ \text{Embedding}({\bm{A}}),\text{in CPCP module}\end{aligned}\right.$
where ${\bm{N}}^{0}$ denotes the initial node feature, ${\bm{P}}$ denotes the PXRD header feature, $t$ represents the time embedding. The directional information between atoms is processed by Fourier transformation as the edge features. Subsequently, the node features, edge features and lattice parameters are all passed through MPNN to obtain final node features and global graph feature. The single-layer structure of MPNN is as follows:
$m_{i,j}^{k}$ $=\text{MLP}({\bm{N}}_{i}^{k},{\bm{N}}_{j}^{k},{\bm{L}}^{\top}{\bm{L}},\text{FT}({\bm{F}}_{i}-{\bm{F}}_{j}))$
$m_{i}^{k}$ $=\sum_{j}m_{i,j}^{k}$
${\bm{N}}_{i}^{k+1}$ $={\bm{N}}_{i}^{k}+\text{MLP}({\bm{N}}_{i}^{k},m_{i}^{k})$
where ${\bm{N}}_{i}^{k}$ denotes the $i$th node feature of ${\bm{N}}^{k}$ in $k$th layer, ${\bm{F}}_{i}$ denotes the $i$th node(atom) coordinate of ${\bm{F}}$ and $\text{FT}(\cdot)$ denotes Fourier transformation. After $s$ layers processing, the global graph feature and the denoising output can be derived as:
${\bm{C}}$ $=\text{Linear}(\frac{1}{N}\sum_{i}^{N}m_{i}^{s})$
$\hat{\varepsilon}_{\bm{F}}^{i}$ $=\text{MLP}({\bm{N}}_{i}^{s})$
$\hat{\varepsilon}_{\bm{L}}$ $={\bm{L}}\cdot\text{MLP}(\frac{1}{N}\sum_{i}^{N}m_{i}^{s})$
where ${\bm{C}}$ denotes the global crystal feature, $\hat{\varepsilon}_{\bm{F}}^{i}$ denotes the $i$th column of the denoising socre $\hat{\varepsilon}_{\bm{F}}$ for fractional coordinates, and $\hat{\varepsilon}_{\bm{L}}$ denotes the denoising term for the lattice. Note that in the CPCP module, ${\bm{P}}$, $t$, $\hat{\varepsilon}_{\bm{F}}$, and $\hat{\varepsilon}_{\bm{L}}$ are not used.
@@
In CPCP module, the crystal feature ${\bm{C}}$ is used as the output of crystal structure network for contrastive learning, which captures the aggregated structural characteristics of the crystal. In CCSG module, $\hat{\varepsilon}_{\bm{F}}$, and $\hat{\varepsilon}_{\bm{L}}$ are the output of crystal struture network. Additionally, ${\bm{P}}$ from the the pre-trained PXRD feature extractor provides conditional PXRD information that guides the denoising direction of the crystal structure. The PXRD header feature serves as a guide throughout reverse process, ensuring the synthesized crystal structure is not only in agreement with the PXRD data but also conforms to the material system’s intrinsic physical and chemical properties. This strategic enhancement integrates critical supervisory data into the model’s architecture.
4.2 Contrastive PXRD-Crystal Pretraining (CPCP) Module
The Contrastive PXRD-Crystal Pretraining (CPCP) module aims to align the embedding spaces of PXRD patterns and crystal structures to facilitate a better understanding of their underlying relationships. In this module, the PXRD feature extractor generates a PXRD header feature vector ${\bm{P}}$ from the PXRD data, and the Crystal Structure Network computes a crystal structure feature vector ${\bm{C}}$ from the crystal graph representation.
To train these embeddings, we use a contrastive learning approach that involves constructing both positive and negative pairs. Positive pairs are formed by pairing PXRD features ${\bm{P}}_{i}$ and crystal structure features ${\bm{C}}_{i}$ that correspond to the same material. Negative pairs, on the other hand, are constructed by pairing a PXRD feature ${\bm{P}}_{i}$ with a crystal structure feature ${\bm{C}}_{j}$ from a different material ($i\neq j$). This setup forces the model to learn robust embeddings by minimizing the distance between positive pairs (similar materials) while maximizing the distance between negative pairs (dissimilar materials).
Constructing negative pairs is crucial as it encourages the model to differentiate between features of different materials, enhancing the model’s discriminative power. The contrastive learning task helps bridge the gap between PXRD data and crystal structures by aligning their representations in a shared embedding space. This alignment allows the model to learn more meaningful correlations between diffraction patterns and their corresponding crystal structures, improving its ability to predict crystal structures from PXRD data.
The loss function used for this task is the InfoNCE loss, which optimizes the embeddings by encouraging high similarity for positive pairs and low similarity for negative pairs. The Contrastive Learning Loss is defined as: